# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 7 Learning-based Robust Motion Planning
7.1 Overview of Existing Motion Planners
@@
    Learning-based motion planner (see, e.g., Rivière et al. [2020], Everett et al. [2018], Berkenkamp et al. [2017], Carius et al. [2020], Ho and Ermon [2016], Gupta et al. [2018], Kuefler et al. [2017]): $(o_{\ell}(x,o_{g}),t)\mapsto u_{d}(o_{g},t)$, modeled by a DNN, where $o_{g}\in\mathbb{R}^{g}$ is a vector containing global environment information, and $o_{\ell}:\mathbb{R}^{n}\times\mathbb{R}^{g}\mapsto\mathbb{R}^{\ell}$ with $\ell\leq g$ is local environment information extracted from $o_{g}\in\mathbb{R}^{g}$ Rivière et al. [2020].
@@
    Robust tube-based motion planner (see, e.g., Singh et al. [2017], Bemporad and Morari [1999], Mayne et al. [2005, 2011], Schulman et al. [2014], Nakka et al. [2021], Manchester and Slotine [2017], Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Sun et al. [2020], Lopez et al. [2019], Lakshmanan et al. [2020], Gahlawat et al. [2021], Sun et al. [2021], Tsukamoto and Chung [2019], Zhao et al. [2021]): $(x,x_{d},u_{d},t)\mapsto u^{*}$, where $u^{*}$ is a robust feedback tracking controller given by Theorems 4.24.66.4, or 6.5.
The robust tube-based motion planner (b) ensures that the perturbed trajectories $x$ of (6.1) and (6.2) stay in an exponentially bounded error tube around the target trajectory $x_{d}$ of (6.3Manchester and Slotine [2017], Singh et al. [2017], Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Sun et al. [2020], Lopez et al. [2019], Lakshmanan et al. [2020], Gahlawat et al. [2021], Sun et al. [2021], Tsukamoto and Chung [2019], Zhao et al. [2021] given as in Fig. 7, using Theorems 2.4 and 2.5. However, it requires the online computation of $(x_{d},u_{d})$ as an input to their control policy given in Theorems 4.24.66.4, and 6.5, which is not realistic for systems with limited computational resources. The learning-based motion planner (a) circumvents this issue by modeling the target policy $(o_{\ell},t)\mapsto u_{d}$ by a DNN. In essence, our approach, to be proposed in Theorem 7.1, is for providing (a) with the contraction theory-based stability guarantees of (b), thereby significantly enhancing the performance of (a) that only assures the tracking error $\|x-x_{d}\|$ to be bounded by a function which increases exponentially with time, as derived previously in Theorem 5.1 of Sec. 5.
@@
7.2 Stability Guarantees of LAG-ROS
The method of LAG-ROS - for Learning-based Autonomous Guidance with RObustness and Stability guarantees Tsukamoto and Chung [2021], bridges the gap between (a) and (b) by ensuring that the distance between the target and controlled trajectories to be exponentially bounded.
@@ 7.2 Stability Guarantees of LAG-ROS
    LAG-ROS (see Fig. 12 and Tsukamoto and Chung [2021]): $(x,o_{\ell}(x,o_{g}),t)\mapsto u^{*}(x,x_{d}(o_{g},t),u_{d}(o_{g},t),t)$, modeled by $u_{L}$ to be given in Theorem 7.1, where $u^{*}(x,x_{d},u_{d},t)$ of (b) is viewed as a function of $(x,o_{\ell},t)$.
@@
Learning-based Autonomous Guidance with RObustness and Stability guarantees (LAG-ROS) is a DNN model for approximating a function which maps state $x$, local environment information $o_{\ell}$, and time $t$ to an optimal robust feedback control input $u^{*}$ given by (3.15) of Theorem 4.2, i.e., $u^{*}=u_{d}-K(x-x_{d})$, or (4.25) of Theorem 4.6, i.e., $u=u_{d}+\int_{0}^{1}kd\mu$, where its training data is sampled as explained in Sec. 7.3 (see Figures 910, and 12).
The LAG-ROS in Definition 7.1 achieves online computation of $u$ without solving a motion planning problem, and it still possesses superior robustness and stability properties due to its internal contracting architecture Tsukamoto and Chung [2021].
@@
Let $u_{L}=u_{L}(x,o_{\ell}(x,o_{g}),t)$ be the LAG-ROS in Definition 7.1, and let $u^{*}$ be given by (3.15) of Theorem 4.2, i.e., $u^{*}=u_{d}-K(x-x_{d})$, or (4.25) of Theorem 4.6, i.e., $u^{*}=u_{d}+\int^{1}_{0}kd\mu$. Define $\Delta_{L}$ of Theorems 5.2 and 5.3 as follows:
@@
where $B=B(x,t)$ is the actuation matrix given in (6.1) and (6.2). Note that we use $q(0,t)=\xi_{0}(t)=x_{d}(o_{g},t)$ and $q(1,t)=\xi_{1}(t)=x(t)$ in the learning-based control formulation of Theorems 5.2 and 5.3.
Table 5: Comparison of LAG-ROS Tsukamoto and Chung [2021] with the Learning-based and Robust Tube-based Planners.
(a) Learning-based $(o_{\ell},t)\mapsto u_{d}$ Increases exponentially (Theorem 5.1) One DNN evaluation
(b) Robust tube-based $(x,x_{d},u_{d},t)\mapsto u^{*}$ Exponentially bounded (Theorems 4.24.66.46.5) Computation of $(x_{d},u_{d})$
(c) LAG-ROS $(x,o_{\ell},t)\mapsto u^{*}$ Exponentially bounded (Theorem 7.1) One DNN evaluation
If the LAG-ROS $u_{L}$ is learned to satisfy (5.3) with $\epsilon_{\ell 1}=0$, i.e., $\|\Delta_{L}\|\leq\epsilon_{\ell 0}$ for all $x\in\mathcal{S}_{s}$, $o_{g}\in\mathcal{S}_{o}$, and $t\in\mathcal{S}_{t}$, where $\mathcal{S}_{s}\subseteq\mathbb{R}^{n}$, $\mathcal{S}_{o}\subseteq\mathbb{R}^{g}$, and $\mathcal{S}_{t}\subseteq\mathbb{R}_{\geq 0}$ are some compact sets, and if the control-affine nonlinear systems (6.1) and (6.2) are controlled by $u=u_{L}$, then (5.18) of Theorem 5.2 and (5.22) of Theorem 5.3 hold with $\bar{d}$ and $\bar{g}$ replaced as in (6.12), i.e., we have the following for $\mathtt{e}(t)=x(t)-x_{d}(o_{g},t)$:
@@
where $x_{d}=x_{d}(o_{g},t)$ and $u_{d}=u_{d}(o_{g},t)$ are as given in (6.3), $\bar{d}_{\epsilon_{\ell}}=\epsilon_{\ell 0}+\bar{d}_{c}$, $C_{\epsilon_{\ell}}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$, and the other variables are as defined in Theorems 5.2 and 5.3 with the disturbance terms in (6.12) of Theorem 6.1.
@@
We have $f(x,t)+B(x,t)u_{L}=f(x,t)+B(x,t)u^{*}+B(x,t)(u_{L}-u^{*})$ and $\|\Delta_{L}\|=\|B(x,t)(u_{L}-u^{*})\|\leq\epsilon_{\ell 0}$ by (5.3) with $\epsilon_{\ell 1}=0$ for $x_{d}=x_{d}(o_{g},t)$ and $u_{d}=u_{d}(o_{g},t)$. Since the virtual system which has $x$ of $\dot{x}=f(x,t)+B(x,t)u^{*}$ and $x_{d}$ of (6.3) as its particular solutions is contracting due to Theorems 4.2 or 4.6, the application of Theorems 5.2 and 5.3 yields the desired relations (7.2)–(7.4). $\qed$
Due to its internal feedback structure, the LAG-ROS achieves exponential boundedness of the trajectory tracking error as in Theorem 7.1, thereby improving the performance of existing learning-based feedforward control laws to model $(o_{\ell},t)\mapsto u_{d}$ as in (a), whose tracking error bound increases exponentially with time as in Theorem 5.1 (see Tsukamoto and Chung [2021]). This property enables its use in safety-critical guidance and control tasks which we often encounter in modern machine learning applications.
@@
Let us consider the cart-pole balancing task in Fig. 8, perturbed externally by deterministic perturbation, to demonstrate the differences of (a)(c) summarized in Table 5. Its dynamical system is given in Tsukamoto et al. [2021], Barto et al. [1983] with its values defined in Barto et al. [1983].
The target trajectories $(x_{d},u_{d})$ are sampled with the objective function $\int_{0}^{T}\|u\|^{2}dt$, and the LAG-ROS of Theorem 7.1 is modeled by a DNN independently of the target trajectory as detailed in Tsukamoto and Chung [2021]. The environment information $o_{g}$ is selected as random initial and target terminal states, and we let $V_{\ell}(0)=0$ and $\epsilon_{\ell 0}+\bar{d}_{c}=0.75$ in (7.2) of Theorem 7.1, yielding the following exponential bound:
@@
Figure 8 shows the tracking errors of each motion planner averaged over $50$ simulations at each time instant $t$. The LAG-ROS (c) and robust tube-based planner (b) indeed satisfy the bound (7.5) (dotted black curve) for all $t$ with a small standard deviation $\sigma$, unlike learning-based motion planner (a) with a diverging bound in Theorem 5.1 and increasing deviation $\sigma$. This example demonstrates one of the major advantages of contraction theory, which enables such quantitative analysis on robustness and stability of learning-based planners.
As implied in Example 7.1, the LAG-ROS indeed possesses the robustness and stability guarantees of $u^{*}$ as proven in Theorem 7.1, unlike (a), while retaining significantly lower computational cost than that of (b). See Tsukamoto and Chung [2021] for a more detailed discussion of this simulation result.
7.3 Tube-based State Constraint Satisfaction
We exploit the result of Theorem 7.1 in the tube-based motion planning Singh et al. [2017] for generating LAG-ROS training data, which satisfies given state constraints even under the existence of the learning error and disturbance of Theorem 5.2. In particular, we sample target trajectories $(x_{d},u_{d})$ of (6.3) in $\mathcal{S}_{s}\times\mathcal{S}_{u}$ of Theorem 7.1 by solving the following, assuming the learning error $\epsilon_{\ell 0}$ and disturbance upper bound $\bar{d}_{c}$ of Theorem 7.1 are selected a priori Shi et al. [2019]:
@@ 7.3 Tube-based State Constraint Satisfaction
where $c_{0}>0$, $c_{1}\geq 0$, $P(\bar{x},\bar{u},t)$ is some performance-based cost function (e.g., information-based cost Nakka et al. [2021]), $\bar{\mathcal{X}}$ is robust admissible state space defined as $\bar{\mathcal{X}}(o_{g},t)=\{v(t)\in\mathbb{R}^{n}|\forall\xi(t)\in\{\xi(t)|\|v(t)-\xi(t)\|\leq r_{\ell}(t)\},~\xi(t)\in\mathcal{X}(o_{g},t)\}$, $\mathcal{X}(o_{g},t)$ is given admissible state space, $r_{\ell}(t)$ is given by (7.2) of Theorem 7.1, and $\bar{u}\in\bar{\mathcal{U}}(o_{g},t)$ is an input constraint. The following theorem shows that the LAG-ROS ensures the perturbed state $x$ of (6.1) to satisfy $x\in\mathcal{X}$, due to the contracting property of $u^{*}$.
@@
If the solution of (7.6) gives $(x_{d},u_{d})$, the LAG-ROS $u_{L}$ of Theorem 7.1 ensures the perturbed solution $x(t)$ of $\dot{x}=f(x,t)+B(x,t)u_{L}+d_{c}(x,t)$ in (6.1) to stay in the admissible state space $\mathcal{X}(o_{g},t)$, i.e., $x(t)\in\mathcal{X}(o_{g},t)$, even with the learning error $\epsilon_{\ell 0}$ of $u_{L}$ in Theorem 7.1.
@@
See Tsukamoto and Chung [2021]. $\qed$
Lemma 7.1 implies that the perturbed trajectory (6.1) controlled by LAG-ROS will not violate the given state constraints as long as $x_{d}$ is sampled by (7.6), which helps greatly reduce the need for safety control schemes such as Kolathaya and Ames [2019]. The localization method in Rivière et al. [2020] allows extracting $o_{\ell}$ of (a) by $o_{g}$ of (7.6), to render LAG-ROS applicable to different environments in a distributed way with a single policy.
We therefore sample artificially perturbed states $x$ in the tube $S(x_{d}(o_{g},t))=\{\xi(t)\in\mathbb{R}^{n}|\|\xi(t)-x_{d}(o_{g},t)\|\leq r_{\ell}(t)\}$ as depicted in Figures 9 and 10 Tsukamoto and Chung [2021]. The robust control inputs $u^{*}$ for training LAG-ROS of Theorem 7.1 is then sampled by computing the CV-STEM of Theorem 4.2 or Theorem 4.6. The LAG-ROS framework is summarized in Fig. 12, and a pseudocode for its offline construction can be found in Tsukamoto and Chung [2021].
@@
For stochastic systems, we could sample $x$ around $x_{d}$ using a given probability distribution of the disturbance using (7.3) and (7.4) of Theorem 7.1.
@@
Let us again consider the spacecraft dynamics of Example 6.2, but now for designing a real-time robust motion planner using the LAG-ROS in a multi-agent setting. In this example, each agent is assumed to have access only to the states of the other agents/obstacles located in a circle with a limited radius, 2.02.0m, centered around it. The performance of LAG-ROS is compared with (a), (b), and the following centralized motion planner (d) which is not computable in real-time:
@@
is used with an input constraint $u_{i}\geq 0,~\forall i$, to avoid collisions with a random number of multiple circular obstacles and of other agents, even under the learning error and disturbances. See Tsukamoto and Chung [2021] for the LAG-ROS training details Tsukamoto and Chung [2021].
@@
  • For (a), the tracking error accumulates exponentially with time due to the lack of robustness as proven in Theorem 5.1.
@@
  • LAG-ROS (c) tackles these two problems by providing formal robustness and stability guarantees of Theorem 7.1, whilst implicitly knowing the global solution (only from the local information $o_{\ell}$ as in Rivière et al. [2020]) without computing it online. It indeed satisfies the state constraints due to Lemma 7.1 as can be seen from Fig. 11.
See Tsukamoto and Chung [2021] for the in-depth discussion on this simulation result.

### v8

@@ 7. Learning-based Robust Motion Planning
7.1. Overview of Existing Motion Planners
@@
    Learning-based motion planner (see, e.g., glas, 8593871, NIPS2017_766ebcd5, 9001182, NIPS2016_cc7e2b87, 8578338, 7995721): $(o_{\ell}(x,o_{g}),t)\mapsto u_{d}(o_{g},t)$, modeled by a DNN, where $o_{g}\in\mathbb{R}^{g}$ is a vector containing global environment information, and $o_{\ell}:\mathbb{R}^{n}\times\mathbb{R}^{g}\mapsto\mathbb{R}^{\ell}$ with $\ell\leq g$ is local environment information extracted from $o_{g}\in\mathbb{R}^{g}$ glas.
@@
    Robust tube-based motion planner (see, e.g., 7989693, 10.1007/BFb0109870, tube_mpc, tube_nmpc, doi:10.1177/0278364914528132, 9290355, ccm, mypaperTAC, ncm, nscm, chuchu, 8814758, 9303957, L1contraction2, sun2021uncertaintyaware, mypaper, zhao2021tubecertified): $(x,x_{d},u_{d},t)\mapsto u^{*}$, where $u^{*}$ is a robust feedback tracking controller given by Theorems LABEL:THM:Thm:CV-STEMLABEL:THM:Thm:ccm_cvstemLABEL:THM:Thm:ncm_clf, or LABEL:THM:Thm:ncm_ccm_clf.
The robust tube-based motion planner (b) ensures that the perturbed trajectories $x$ of (6.1) and (6.2) stay in an exponentially bounded error tube around the target trajectory $x_{d}$ of (6.3ccm, 7989693, mypaperTAC, ncm, nscm, chuchu, 8814758, 9303957, L1contraction2, sun2021uncertaintyaware, mypaper, zhao2021tubecertified given as in Fig. 7, using Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic. However, it requires the online computation of $(x_{d},u_{d})$ as an input to their control policy given in Theorems LABEL:THM:Thm:CV-STEMLABEL:THM:Thm:ccm_cvstemLABEL:THM:Thm:ncm_clf, and LABEL:THM:Thm:ncm_ccm_clf, which is not realistic for systems with limited computational resources. The learning-based motion planner (a) circumvents this issue by modeling the target policy $(o_{\ell},t)\mapsto u_{d}$ by a DNN. In essence, our approach, to be proposed in Theorem LABEL:THM:Thm:lagros_stability, is for providing (a) with the contraction theory-based stability guarantees of (b), thereby significantly enhancing the performance of (a) that only assures the tracking error $\|x-x_{d}\|$ to be bounded by a function which increases exponentially with time, as derived previously in Theorem LABEL:THM:Thm:naive_learning of Sec. 5.
@@
7.2. Stability Guarantees of LAG-ROS
The method of LAG-ROS - for Learning-based Autonomous Guidance with RObustness and Stability guarantees lagros, bridges the gap between (a) and (b) by ensuring that the distance between the target and controlled trajectories to be exponentially bounded.
@@ 7.2. Stability Guarantees of LAG-ROS
    LAG-ROS (see Fig. 12 and lagros): $(x,o_{\ell}(x,o_{g}),t)\mapsto u^{*}(x,x_{d}(o_{g},t),u_{d}(o_{g},t),t)$, modeled by $u_{L}$ to be given in Theorem LABEL:THM:Thm:lagros_stability, where $u^{*}(x,x_{d},u_{d},t)$ of (b) is viewed as a function of $(x,o_{\ell},t)$.
@@
LAG-ROSDef:lagros  Learning-based Autonomous Guidance with RObustness and Stability guarantees (LAG-ROS) is a DNN model for approximating a function which maps state $x$, local environment information $o_{\ell}$, and time $t$ to an optimal robust feedback control input $u^{*}$ given by (3.15) of Theorem LABEL:THM:Thm:CV-STEM, i.e., $u^{*}=u_{d}-K(x-x_{d})$, or (4.25) of Theorem LABEL:THM:Thm:ccm_cvstem, i.e., $u=u_{d}+\int_{0}^{1}kd\mu$, where its training data is sampled as explained in Sec. 7.3 (see Figures 910, and 12).
The LAG-ROS in Definition LABEL:DEF:Def:lagros achieves online computation of $u$ without solving a motion planning problem, and it still possesses superior robustness and stability properties due to its internal contracting architecture lagros.
@@
Learning-based Motion Planning with Contraction TheoryThm:lagros_stability Let $u_{L}=u_{L}(x,o_{\ell}(x,o_{g}),t)$ be the LAG-ROS in Definition LABEL:DEF:Def:lagros, and let $u^{*}$ be given by (3.15) of Theorem LABEL:THM:Thm:CV-STEM, i.e., $u^{*}=u_{d}-K(x-x_{d})$, or (4.25) of Theorem LABEL:THM:Thm:ccm_cvstem, i.e., $u^{*}=u_{d}+\int^{1}_{0}kd\mu$. Define $\Delta_{L}$ of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto as follows:
@@
where $B=B(x,t)$ is the actuation matrix given in (6.1) and (6.2). Note that we use $q(0,t)=\xi_{0}(t)=x_{d}(o_{g},t)$ and $q(1,t)=\xi_{1}(t)=x(t)$ in the learning-based control formulation of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto.
If the LAG-ROS $u_{L}$ is learned to satisfy (5.3) with $\epsilon_{\ell 1}=0$, i.e., $\|\Delta_{L}\|\leq\epsilon_{\ell 0}$ for all $x\in\mathcal{S}_{s}$, $o_{g}\in\mathcal{S}_{o}$, and $t\in\mathcal{S}_{t}$, where $\mathcal{S}_{s}\subseteq\mathbb{R}^{n}$, $\mathcal{S}_{o}\subseteq\mathbb{R}^{g}$, and $\mathcal{S}_{t}\subseteq\mathbb{R}_{\geq 0}$ are some compact sets, and if the control-affine nonlinear systems (6.1) and (6.2) are controlled by $u=u_{L}$, then (5.18) of Theorem LABEL:THM:Thm:contraction_learning and (5.22) of Theorem LABEL:THM:Thm:contraction_learning_sto hold with $\bar{d}$ and $\bar{g}$ replaced as in (6.12), i.e., we have the following for $\mathtt{e}(t)=x(t)-x_{d}(o_{g},t)$:
@@
where $x_{d}=x_{d}(o_{g},t)$ and $u_{d}=u_{d}(o_{g},t)$ are as given in (6.3), $\bar{d}_{\epsilon_{\ell}}=\epsilon_{\ell 0}+\bar{d}_{c}$, $C_{\epsilon_{\ell}}=\bar{g}_{c}^{2}({2}{\alpha_{G}}^{-1}+1)+\epsilon_{\ell 0}^{2}\alpha_{d}^{-1}$, and the other variables are as defined in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto with the disturbance terms in (6.12) of Theorem LABEL:THM:Thm:NCMstability1.
@@
We have $f(x,t)+B(x,t)u_{L}=f(x,t)+B(x,t)u^{*}+B(x,t)(u_{L}-u^{*})$ and $\|\Delta_{L}\|=\|B(x,t)(u_{L}-u^{*})\|\leq\epsilon_{\ell 0}$ by (5.3) with $\epsilon_{\ell 1}=0$ for $x_{d}=x_{d}(o_{g},t)$ and $u_{d}=u_{d}(o_{g},t)$. Since the virtual system which has $x$ of $\dot{x}=f(x,t)+B(x,t)u^{*}$ and $x_{d}$ of (6.3) as its particular solutions is contracting due to Theorems LABEL:THM:Thm:CV-STEM or LABEL:THM:Thm:ccm_cvstem, the application of Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto yields the desired relations (7.2)–(7.4). $\qed$
Table 5: Comparison of LAG-ROS lagros with the Learning-based and Robust Tube-based Planners.
(a) Learning-based $(o_{\ell},t)\mapsto u_{d}$ Increases exponentially (Theorem LABEL:THM:Thm:naive_learning) One DNN evaluation
(b) Robust tube-based $(x,x_{d},u_{d},t)\mapsto u^{*}$ Exponentially bounded (Theorems LABEL:THM:Thm:CV-STEMLABEL:THM:Thm:ccm_cvstemLABEL:THM:Thm:ncm_clfLABEL:THM:Thm:ncm_ccm_clf) Computation of $(x_{d},u_{d})$
(c) LAG-ROS $(x,o_{\ell},t)\mapsto u^{*}$ Exponentially bounded (Theorem LABEL:THM:Thm:lagros_stability) One DNN evaluation
Due to its internal feedback structure, the LAG-ROS achieves exponential boundedness of the trajectory tracking error as in Theorem LABEL:THM:Thm:lagros_stability, thereby improving the performance of existing learning-based feedforward control laws to model $(o_{\ell},t)\mapsto u_{d}$ as in (a), whose tracking error bound increases exponentially with time as in Theorem LABEL:THM:Thm:naive_learning (see lagros). This property enables its use in safety-critical guidance and control tasks which we often encounter in modern machine learning applications.
@@
Cart-Pole Balancingex:cartpole Let us consider the cart-pole balancing task in Fig. 8, perturbed externally by deterministic perturbation, to demonstrate the differences of (a)(c) summarized in Table 5. Its dynamical system is given in ancm, 6313077 with its values defined in 6313077.
The target trajectories $(x_{d},u_{d})$ are sampled with the objective function $\int_{0}^{T}\|u\|^{2}dt$, and the LAG-ROS of Theorem LABEL:THM:Thm:lagros_stability is modeled by a DNN independently of the target trajectory as detailed in lagros. The environment information $o_{g}$ is selected as random initial and target terminal states, and we let $V_{\ell}(0)=0$ and $\epsilon_{\ell 0}+\bar{d}_{c}=0.75$ in (7.2) of Theorem LABEL:THM:Thm:lagros_stability, yielding the following exponential bound:
@@
Figure 8 shows the tracking errors of each motion planner averaged over $50$ simulations at each time instant $t$. The LAG-ROS (c) and robust tube-based planner (b) indeed satisfy the bound (7.5) (dotted black curve) for all $t$ with a small standard deviation $\sigma$, unlike learning-based motion planner (a) with a diverging bound in Theorem LABEL:THM:Thm:naive_learning and increasing deviation $\sigma$. This example demonstrates one of the major advantages of contraction theory, which enables such quantitative analysis on robustness and stability of learning-based planners.
As implied in Example LABEL:EX:ex:cartpole, the LAG-ROS indeed possesses the robustness and stability guarantees of $u^{*}$ as proven in Theorem LABEL:THM:Thm:lagros_stability, unlike (a), while retaining significantly lower computational cost than that of (b). See lagros for a more detailed discussion of this simulation result.
7.3. Tube-based State Constraint Satisfaction
We exploit the result of Theorem LABEL:THM:Thm:lagros_stability in the tube-based motion planning 7989693 for generating LAG-ROS training data, which satisfies given state constraints even under the existence of the learning error and disturbance of Theorem LABEL:THM:Thm:contraction_learning. In particular, we sample target trajectories $(x_{d},u_{d})$ of (6.3) in $\mathcal{S}_{s}\times\mathcal{S}_{u}$ of Theorem LABEL:THM:Thm:lagros_stability by solving the following, assuming the learning error $\epsilon_{\ell 0}$ and disturbance upper bound $\bar{d}_{c}$ of Theorem LABEL:THM:Thm:lagros_stability are selected a priori neurallander:
@@ 7.3. Tube-based State Constraint Satisfaction
where $c_{0}>0$, $c_{1}\geq 0$, $P(\bar{x},\bar{u},t)$ is some performance-based cost function (e.g., information-based cost doi:10.2514/6.2021-1103), $\bar{\mathcal{X}}$ is robust admissible state space defined as $\bar{\mathcal{X}}(o_{g},t)=\{v(t)\in\mathbb{R}^{n}|\forall\xi(t)\in\{\xi(t)|\|v(t)-\xi(t)\|\leq r_{\ell}(t)\},~\xi(t)\in\mathcal{X}(o_{g},t)\}$, $\mathcal{X}(o_{g},t)$ is given admissible state space, $r_{\ell}(t)$ is given by (7.2) of Theorem LABEL:THM:Thm:lagros_stability, and $\bar{u}\in\bar{\mathcal{U}}(o_{g},t)$ is an input constraint. The following theorem shows that the LAG-ROS ensures the perturbed state $x$ of (6.1) to satisfy $x\in\mathcal{X}$, due to the contracting property of $u^{*}$.
@@
If the solution of (7.6) gives $(x_{d},u_{d})$, the LAG-ROS $u_{L}$ of Theorem LABEL:THM:Thm:lagros_stability ensures the perturbed solution $x(t)$ of $\dot{x}=f(x,t)+B(x,t)u_{L}+d_{c}(x,t)$ in (6.1) to stay in the admissible state space $\mathcal{X}(o_{g},t)$, i.e., $x(t)\in\mathcal{X}(o_{g},t)$, even with the learning error $\epsilon_{\ell 0}$ of $u_{L}$ in Theorem LABEL:THM:Thm:lagros_stability.
@@
See lagros. $\qed$
Lemma 7.1 implies that the perturbed trajectory (6.1) controlled by LAG-ROS will not violate the given state constraints as long as $x_{d}$ is sampled by (7.6), which helps greatly reduce the need for safety control schemes such as 8405547. The localization method in glas allows extracting $o_{\ell}$ of (a) by $o_{g}$ of (7.6), to render LAG-ROS applicable to different environments in a distributed way with a single policy.
We therefore sample artificially perturbed states $x$ in the tube $S(x_{d}(o_{g},t))=\{\xi(t)\in\mathbb{R}^{n}|\|\xi(t)-x_{d}(o_{g},t)\|\leq r_{\ell}(t)\}$ as depicted in Figures 9 and 10 lagros. The robust control inputs $u^{*}$ for training LAG-ROS of Theorem LABEL:THM:Thm:lagros_stability is then sampled by computing the CV-STEM of Theorem LABEL:THM:Thm:CV-STEM or Theorem LABEL:THM:Thm:ccm_cvstem. The LAG-ROS framework is summarized in Fig. 12, and a pseudocode for its offline construction can be found in lagros.
@@
For stochastic systems, we could sample $x$ around $x_{d}$ using a given probability distribution of the disturbance using (7.3) and (7.4) of Theorem LABEL:THM:Thm:lagros_stability.
@@
Planar Spacecraft Simulatorsex:sc_lagros Let us again consider the spacecraft dynamics of Example LABEL:EX:ex:sc_robust_control, but now for designing a real-time robust motion planner using the LAG-ROS in a multi-agent setting. In this example, each agent is assumed to have access only to the states of the other agents/obstacles located in a circle with a limited radius, 2.02.0m, centered around it. The performance of LAG-ROS is compared with (a), (b), and the following centralized motion planner (d) which is not computable in real-time:
@@
is used with an input constraint $u_{i}\geq 0,~\forall i$, to avoid collisions with a random number of multiple circular obstacles and of other agents, even under the learning error and disturbances. See lagros for the LAG-ROS training details lagros.
@@
  • For (a), the tracking error accumulates exponentially with time due to the lack of robustness as proven in Theorem LABEL:THM:Thm:naive_learning.
@@
  • LAG-ROS (c) tackles these two problems by providing formal robustness and stability guarantees of Theorem LABEL:THM:Thm:lagros_stability, whilst implicitly knowing the global solution (only from the local information $o_{\ell}$ as in glas) without computing it online. It indeed satisfies the state constraints due to Lemma 7.1 as can be seen from Fig. 11.
See lagros for the in-depth discussion on this simulation result.